While driving on the freeway at 110 km/h, you pass a
truck whose total length you estimate at 25 m. (a) If it takes you, in the driver’s seat, 5.5 s to pass from the rear of the truck to its front, what is the truck’s speed relative to the road? (b) How far does the truck travel while you’re passing it?
1
Expert's answer
2018-10-09T07:20:09-0400
a) The relative speed is
v-u=l/t
The truck’s speed relative to the road is
u=v-l/t=110/3.6-25/5.5=26 m/s.
b)
d=ut=(26)(5.5)=143 m.